OpenStudy (i0iz0gangster):

any help would be awesome :) simplify (8+ sqrt 11)(8- sqrt 11) simplify 2 sqrt 6 + 3 sqrt 96 simplify ( 8+ sqrt 11)(8-sqrt 11)

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] and so \[(8+\sqrt{11})(8-\sqrt{11})=8^2-11\]

OpenStudy (i0iz0gangster):

quick question before I say thank you @satellite73 did you do the equations with in the parentheses then multiplied ?

OpenStudy (jazzyfa30):

yes

OpenStudy (i0iz0gangster):

do you mind explaining how to solve the second one

OpenStudy (jazzyfa30):

ok

OpenStudy (jazzyfa30):

\[2 \sqrt 6 + 3 \sqrt 96\]

OpenStudy (jazzyfa30):

ummmmmm hold on

OpenStudy (i0iz0gangster):

nvm lol I got it

OpenStudy (i0iz0gangster):

thank u sooooo much

OpenStudy (jazzyfa30):

for what